# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

### v3

@@ Appendix E Evaluation Results of Description Task based on Traditional Image Captioning Metrics
Table 7 shows the performance of models on traditional image captioning evaluation metrics. Following Krause et al. [26], we use METEOR [35], CIDEr [36], BLEU-1, BLEU-2, BLEU-3, and BLEU-4 [37] to evaluate the performance of models on Description task. Similar to the findings of Zhu et al. [18], it can be observed that traditional image captioning evaluation metrics are not quite suitable for evaluating the Description task. There are two possible reasons. The first possible reason is that image descriptions are longer and more flexible than traditional image captions. The second possible reason is that Description task requires evalution metrics to consider high-level semantics in description, while traditional image captioning evaluation metrics only concerns low-level information.
@@
Appendix F Implementation of Non-GPT-based Cognition Evaluation Methods of Description Task
@@
The task of cognition evaluation of Description task is to determine whether a CoR is mentioned in the description. Apart from ChatGPT or GPT-4, some other evaluation methods are implemented to perform this classification task, as shown in Table 4.
@@
For methods based on ROUGE [38], BERTScore [39], and BLEURT [40], we first split the description into sentences, then use a CoR as a reference to calculate the (recall) score for each sentence compared to the CoR. Then, the highest score among all calculated scores is taken as the score of the CoR corresponding to the description. Finally, the score is converted into 0/1 using a threshold.
@@
We also tried Natural Language Inference (NLI) models to perform the task. First, we use DeBERTa [41] to perform sentence-level NLI task similar to methods mentioned above. If there is at least one “Entailment” for all the sentences, the score of the CoR will be 1. The model we adopted is mDeBERTa-v3-base-xnli-multilingual-nli-2mil7. The second NLI model we tried is DocNLI [42], which can directly take the description and CoR as input and do the classification task.
Appendix G Ethical Considerations
Most images in CogBench are manually collected from Pinterest. We follow the terms of service of Pinterest to collect these images. The images are used as fair use for research purposes only. We will share the data with other researchers who will follow the ethical considerations as established in this study. We have compensated our annotators at a rate of \$0.60 per image, which aligns with the university’s recruitment standards for labeling personnel. This amount is offered as an appropriate token of gratitude for their valuable time and efforts in assisting with our project. We answer any questions our annotators have during their annotation process promptly. They are free to take breaks or quit the annotation task at any time.

### v4

@@ Appendix E Evaluation of LVLMs on the Description Task Using Traditional Image Captioning Metrics
Table 7 shows the model performance on traditional image captioning evaluation metrics. Following Krause et al. (2017), we use METEOR Banerjee and Lavie (2005), CIDEr Vedantam et al. (2015), BLEU-1, BLEU-2, BLEU-3, and BLEU-4 Papineni et al. (2002) to evaluate the model performance on the Description task. Similar to the findings of Zhu et al. (2023a), it can be observed that traditional image captioning evaluation metrics are not quite suitable for evaluating the Description task. There are two possible reasons. The first possible reason is that image descriptions are longer and more flexible than traditional image captions. The second possible reason is that our Description task requires evaluation metrics to place more emphasis on high-level semantics in the description.
@@
Appendix F Implementation of Non-GPT-Based Cognition Evaluation Methods for the Description Task
@@
The cognition evaluation of the Description task is to determine whether each CoR is mentioned in the description. Apart from ChatGPT or GPT-4, some other evaluation methods are also implemented to perform this classification task, as shown in Table 4.
@@
For methods based on ROUGE Lin (2004), BERTScore Zhang et al. (2020), and BLEURT Sellam et al. (2020), we first split the description into sentences and then use each CoR as the reference to calculate the (recall) score for each sentence compared to the CoR. Then, the highest score among all calculated scores is taken as the score of the CoR corresponding to the description. Finally, the score is converted into 0 or 1 using a threshold.
@@
We also tried Natural Language Inference (NLI) models to perform the task. First, we use DeBERTa He et al. (2021, 2023) to perform a sentence-level NLI task similar to the method mentioned above. If there is at least one “Entailment” for all the sentences, the score of the CoR will be 1. The model we adopt is mDeBERTa-v3-base-xnli-multilingual-nli-2mil7. The second NLI model we tried is DocNLI Yin et al. (2021), which can directly take the description and CoR as input and perform the classification task.